OpenStudy (anonymous):

34.8 mi/hr converted to km/min

OpenStudy (anonymous):

\[34.8 \frac{mi}{hr}(\frac{1.602 Km}{1 mi})(\frac{1 hr}{60 \min})=the answer\]

OpenStudy (anonymous):

use the factor label above but change 1.602 to 1.609